# Selfless reduced free product $C^{*}$-algebras

### v1

@@ 3.2. Tracial polynomial growth
(See [54]). If $A$ admits a filtration that satisfies Voiculescu’s condition, then $A$ is nuclear, by [33, Corollary 2.2], and by [60, Corollary 5.4] every quotient of $A$ has a tracial state, i.e., $A$ has the QTS property. Thus, if $A$ has tracial polynomial growth, then $A$ has rapid decay, it is nuclear, and has the QTS property. We note that, by [40, Theorem 2], a $C^{*}$-algebra is nuclear and with the QTS property if and only if it is symmetrically amenable.
@@
Using that $k_{n}\lesssim(n+1)^{\gamma}$ completes the proof.
@@ 3.3. Examples
Suppose that $a<b$ are real numbers and that $\mu$ is an atomless Borel probability measure whose support is the interval $[a,b]$. Then $(C([a,b],\int\cdot\,d\mu)$ is isomorphic, as a $C^{*}$-probability space, to $(C([0,1]),\int\cdot\,dm)$ (where $m$ denotes the Lebesgue measure). Indeed, since $\mu$ is atomless and fully supported on $[a,b]$, $F\colon[a,b]\to[0,1]$ given by $F(t)=\mu([a,t])$ is a homeomorphism such that $F_{*}(\mu)$ is the Lebesgue measure. Thus, we can always endow $C([a,b])$ with a filtration having tracial polynomial growth (and the rapid decay property) through its identification with $(C([0,1]),\int\cdot\,dm)$ (see previous example).
@@
That is, $E_{\pi}$ is the span of the ‘‘matrix coefficients’’ of $\pi$. Let $\widehat{G}$ denote the set of irreducible representations of $G$. It is shown in [56, Example 3.5], and in more detail in [6, Theorem 2.1], that a length function $\ell\colon\widehat{G}\to\{0,1,\ldots\}$ can be defined such that the spaces
@@ 4.2. Proof of the rapid decay property for free products
(4.4) $\left(\sum_{(I,\widetilde{I})\in{\mathcal{B}}_{r-1}\times{\mathcal{B}}_{\ell-r}}\|T_{I,\widetilde{I}}\|^{2}\right)^{1/2}\leq\sqrt{m}\left(\max_{j}C_{j}\right)\left(\sum_{I\in{\mathcal{B}}_{\ell}}|\lambda_{\widehat{I}}|^{2}\right)^{1/2}=\sqrt{m}(t+1)^{\max_{j}(\alpha_{j})}\|y\|_{2}.$
@@ 5. Selflessness via an Avitzour condition
Fix $m\in{\mathbb{N}}$ with $m>k$ and fix $n>m$. We wish to understand how far in the filtration $(E_{s})_{s}$ the subspace $\phi_{n}(V_{n})$ is. Suppose that $a\in B_{1}*_{\textnormal{alg}}B_{2}*_{\textnormal{alg}}B_{3}$ is an alternating centered word in $(\iota_{1}(V_{n,1}),\iota_{2}(V_{n,2},\iota_{1}(V_{n,3}))$ of length $\ell>0$ with $\ell\leq m$. As in the proof of Lemma 5.3, write
@@ 6. Selflessness for more general free products
The main idea of the proof of Theorem 5.1 can be adapted to show that $(A,\rho)*(C({\mathbb{T}}),m)$ is selfless, where $m$ is integration against Lebesgue measure and $(A,\rho)$ has a filtration with the rapid decay property. Indeed in this case, we simply adapt Lemma 5.3 using the *-homomorphism $\phi_{n}\colon B_{1}*_{\textnormal{alg}}B_{2}*_{\textnormal{alg}}B_{3}\to A*C({\mathbb{T}})$ (where $B_{1}=B_{3}=A,B_{2}=C({\mathbb{T}})$), defined by declaring that $\phi_{n}$ is the identity on $B_{1},B_{2}$ and is conjugation by $u^{n}$ on $B_{3}$, where $u$ is the Haar unitary generating $C({\mathbb{T}})$. The proof precedes, mutatis mutandis, as that of Theorem 5.3. In this section, we expand on, and abstract, this strategy for free products with the circle to handle free products $(A_{1},\rho)*(A_{2},\tau)$ where $\tau$ is a trace.
@@ 6.1. The general approach
Note that $y_{0},y_{t}$ could be empty in which case $y_{0}v=v$ and $v^{*}y_{t}=v^{*}$. If $y_{0}$ is not empty, then it ends in $(\Upsilon_{2}\cup\{1\})v$ and other letters are in $\Upsilon_{1}\cup\Upsilon_{2}$. Similarly, if $y_{t}$ is not empty then it is an alternating word in $(\Upsilon_{1},\Upsilon_{2}\cup(\Upsilon_{2}\cup\{1\})v^{*})$. If $11\leq i\leq t-$ 1 and $y_{i}$ is a single letter in $\Upsilon_{2}$, then $v^{*}y_{i}v$ is a single letter in $v^{*}\Upsilon_{2}v$. If $y_{i}$ is a single letter in $\Upsilon_{1}$, then $v^{*}y_{i}v$ is an alternating word in $(\Upsilon_{1},\{v,v^{*}\})$. Finally, if $y_{i}$ has length longer than $2$, then $v^{*}y_{i}v$ is an alternating word in
@@
To finish the proof of the first item, note that by assumption we can find $b_{i,k}\in F_{k}$ such that $\|\widehat{b}_{i,k}-b_{i,k}\|\to_{k\to\infty}0$. Set $b_{k}=b_{1,k}\cdots b_{p,k}$ and observe that
@@ 7. Free products of separable $C^{*}$-algebras with rapid decay
The goal of this section is to use Theorem 6.4 to prove the following.
@@
Let $M$ be the GNS completion of $A$ with respect to $\tau$. We continue to use $\tau$ for the trace on $M$. Fix an ultrafilter $\omega\in\beta{\mathbb{N}}\setminus{\mathbb{N}}$ and let $({\mathcal{M}},\tau_{\omega})$ be the tracial ultrapower of $(M,\tau)$. By separability of $A$, we can apply [44] to find a Haar unitary $u\in{\mathcal{U}}({\mathcal{M}})$ which is free from $M$. Write $u=e^{ih}$ with $h=h^{*}\in{\mathcal{M}}$. By [32, Theorem 3.3] we may lift $a$ to a uniformly bounded sequence $(h_{k})_{k}\in A_{s.a.}$ Set $u_{k}=e^{ih_{k}}$, such that $u_{k}\in{\mathcal{U}}(A)$. Since continuous functional calculus commutes with *-homomorphisms, we have that $(u_{k})_{k}$ is a lift of $u.$ Thus:
@@ 8. Selflessness for free products of finite dimensional abelian $C^{*}$-algebras
We derive this theorem from Theorem 7.1 from the previous section combined with Dykema’s results on the structure of reduced free products of finite dimensional abelian $C^{*}$-algebras.
@@
Let $p_{2}$ be the unit of $A_{2}$. Since $A$ is simple, $p_{2}$ is a full projection, and so $A$ is stably isomorphic to $p_{2}Ap_{2}$. In the tracial case, selflessness is preserved under stable isomorphism, by [47, Theorems 4.3, 4.4]. It thus suffices to show that $(p_{2}Ap_{2},\frac{1}{\tau(p_{2})}\tau)$ is selfless. By [17, Proposition 2.8],
@@
The conditions (8.1) then read as $\alpha_{i}+\beta_{j}<1$ for all $i,j$. By [19, Theorem 1], these conditions, together with $m+n\geq 5$, are necessary and sufficient for the reduced free product $C^{*}$-algebra $C$ to be a simple $C^{*}$-algebra. They are thus also necessary for $(C,\tau)$ to be selfless.
Let us prove sufficiency. Suppose that $m+n\geq 5$ and that $\alpha_{i}+\beta_{j}<1$ for all $i,j$. Assume without loss of generality that $m\geq 3$ and $n\geq 2$. Let $p=p_{1}+p_{2}$. Since $C$ is a simple $C^{*}$-algebra (by[19, Theorem 1]), $pCp$ is stably isomorphic to $A$. By the permanence of the selfless property under stable isomorphism in the tracial case ([47, Theorems 4.3, 4.4]), it suffices to show that $(pCp,\frac{1}{\tau(p)}\tau)$ is selfless. By [17, Proposition 2.8], we have that
@@
Let us consider first the case when $n+m>5$. In this case, by Dykema’s structure theorem on reduced free products of finite abelian $C^{*}$-algebras [17, Theorem 1],
@@
If $\alpha_{1}+\alpha_{2}+\beta_{j}=1$ for some $j$, then $\alpha_{3}=\beta_{j}$. But this implies that $\alpha_{3}+(1-\beta_{j})=1$, which violates (8.1). Thus, $\alpha_{1}+\alpha_{2}+\beta_{j}\neq 1$ for $j=0,1$. It follows by the structure theorem for reduced free products of the form ${\mathbb{C}}^{2}*{\mathbb{C}}^{2}$ [17, Proposition 2.7], that

### v2

@@ 3.2. Tracial polynomial growth
(See [59]). If $A$ admits a filtration that satisfies Voiculescu’s condition, then $A$ is nuclear, by [36, Corollary 2.2], and by [65, Corollary 5.4] every quotient of $A$ has a tracial state, i.e., $A$ has the QTS property. Thus, if $A$ has tracial polynomial growth, then $A$ has rapid decay, it is nuclear, and has the QTS property. We note that, by [43, Theorem 2], a $C^{*}$-algebra is nuclear and with the QTS property if and only if it is symmetrically amenable.
@@
Using that $k_{n}\lesssim(n+1)^{\gamma}$ completes the proof. ∎
@@ 3.3. Examples
Suppose that $a<b$ are real numbers and that $\mu$ is an atomless Borel probability measure whose support is the interval $[a,b]$. Then $(C([a,b]),\int\cdot\,d\mu)$ is isomorphic, as a $C^{*}$-probability space, to $(C([0,1]),\int\cdot\,dm)$ (where $m$ denotes the Lebesgue measure). Indeed, since $\mu$ is atomless and fully supported on $[a,b]$, $F\colon[a,b]\to[0,1]$ given by $F(t)=\mu([a,t])$ is a homeomorphism such that $F_{*}(\mu)$ is the Lebesgue measure. Thus, we can always endow $C([a,b])$ with a filtration having tracial polynomial growth (and the rapid decay property) through its identification with $(C([0,1]),\int\cdot\,dm)$ (see previous example).
@@
That is, $E_{\pi}$ is the span of the ‘‘matrix coefficients’’ of $\pi$. Let $\widehat{G}$ denote the set of irreducible representations of $G$. It is shown in [61, Example 3.5], and in more detail in [6, Theorem 2.1], that a length function $\ell\colon\widehat{G}\to\{0,1,\ldots\}$ can be defined such that the spaces
@@ 4.2. Proof of the rapid decay property for free products
(4.4) $\left(\sum_{(I,\widetilde{I})\in{\mathcal{B}}_{r-1}\times{\mathcal{B}}_{\ell-r}}\|T_{I,\widetilde{I}}\|^{2}\right)^{1/2}$ $\leq\sqrt{m}\left(\max_{j}C_{j}\right)\left(\sum_{I\in{\mathcal{B}}_{\ell}}|\lambda_{\widehat{I}}|^{2}\right)^{1/2}$
$=\sqrt{m}\left(\max_{j}C_{j}\right)\|y\|_{2}.$
@@ 5. Selflessness via an Avitzour condition
Fix $m\in{\mathbb{N}}$ with $m>k$ and fix $n>m$. We wish to understand how far in the filtration $(E_{s})_{s}$ the subspace $\phi_{n}(V_{n})$ is. Suppose that $a\in B_{1}*_{\textnormal{alg}}B_{2}*_{\textnormal{alg}}B_{3}$ is an alternating centered word in $(\iota_{1}(V_{n,1}),\iota_{2}(V_{n,2}),\iota_{3}(V_{n,1}))$ of length $\ell>0$ with $\ell\leq m$. As in the proof of Lemma 5.3, write
@@ 6. Selflessness for more general free products
The main idea of the proof of Theorem 5.1 can be adapted to show that $(A,\rho)*(C({\mathbb{T}}),m)$ is selfless, where $m$ is integration against Lebesgue measure and $(A,\rho)$ has a filtration with the rapid decay property. Indeed in this case, we simply adapt Lemma 5.3 using the *-homomorphism $\phi_{n}\colon B_{1}*_{\textnormal{alg}}B_{2}*_{\textnormal{alg}}B_{3}\to A*C({\mathbb{T}})$ (where $B_{1}=B_{3}=A,B_{2}=C({\mathbb{T}})$), defined by declaring that $\phi_{n}$ is the identity on $B_{1},B_{2}$ and is conjugation by $u^{n}$ on $B_{3}$, where $u$ is the Haar unitary generating $C({\mathbb{T}})$. The proof proceeds, with minor modifications, as that of Theorem 5.3. In this section, we expand on, and abstract, this strategy for free products with the circle to handle free products $(A_{1},\rho)*(A_{2},\tau)$ where $\tau$ is a trace.
@@ 6.1. The general approach
Note that $y_{0},y_{t}$ could be empty in which case $y_{0}v=v$ and $v^{*}y_{t}=v^{*}$. If $y_{0}$ is not empty, then $y_{0}v$ ends in $(\Upsilon_{2}\cup\{1\})v$ and its other letters are in $\Upsilon_{1}\cup\Upsilon_{2}$. Similarly, if $y_{t}$ is not empty, then $v^{*}y_{t}$ is an alternating word in $(\Upsilon_{1},\Upsilon_{2}\cup v^{*}(\Upsilon_{2}\cup\{1\}))$. If $11\leq i\leq t-$ 1 and $y_{i}$ is a single letter in $\Upsilon_{2}$, then $v^{*}y_{i}v$ is a single letter in $v^{*}\Upsilon_{2}v$. If $y_{i}$ is a single letter in $\Upsilon_{1}$, then $v^{*}y_{i}v$ is an alternating word in $(\Upsilon_{1},\{v,v^{*}\})$. Finally, if $y_{i}$ has length longer than $2$, then $v^{*}y_{i}v$ is an alternating word in
@@
To finish the proof of the first item, note that by assumption we can find $b_{i,k}\in\widehat{F}_{k}$ such that $\|\widehat{b}_{i,k}-b_{i,k}\|\to_{k\to\infty}0$. Set $b_{k}=b_{1,k}\cdots b_{p,k}$ and observe that
@@ 7. Free products of separable $C^{*}$-algebras with rapid decay
The goal of this section is to use Theorem 6.4 to prove Theorem B from the introduction, which we restate here:
@@
Let $M$ be the GNS completion of $A$ with respect to $\tau$. We continue to use $\tau$ for the trace on $M$. Fix an ultrafilter $\omega\in\beta{\mathbb{N}}\setminus{\mathbb{N}}$ and let $({\mathcal{M}},\tau_{\omega})$ be the tracial ultrapower of $(M,\tau)$. By separability of $A$, we can apply [47] to find a Haar unitary $u\in{\mathcal{U}}({\mathcal{M}})$ which is free from $M$. Write $u=e^{ih}$ with $h=h^{*}\in{\mathcal{M}}$. By [35, Theorem 3.3] we may lift $a$ to a uniformly bounded sequence $(h_{k})_{k}\in A_{s.a.}$ Set $u_{k}=e^{ih_{k}}$, such that $u_{k}\in{\mathcal{U}}(A)$. Since continuous functional calculus commutes with *-homomorphisms, we have that $(u_{k})_{k}$ is a lift of $u.$ Thus:
@@
Let $H_{\mathbb{R}}$ be a real Hilbert space and $(U_{t})_{t}$ a one-parameter group of orthogonal transformations on $H_{\mathbb{R}}$. Let $\Gamma(H_{\mathbb{R}},U_{t})$ denote the free Araki-Wood $C^{*}$-algebra introduced by Shlyakhtenko in [55]. Let $\phi_{U}$ denote the vacuum state on $\Gamma(H_{\mathbb{R}},U_{t})$. We refer the reader to [55] for details on this construction. We note that, in the case that $(U_{t})_{t}$ acts trivially on $H_{\mathbb{R}}$, $\Gamma(H_{\mathbb{R}},U_{t})$ agrees with the $C^{*}$-algebra generated by a free semicircular system. In the case that $(U_{t})_{t}$ is nontrivial, Shlyakhtenko shows that $\Gamma(H_{\mathbb{R}},U_{t})$ is simple and traceless.
Corollary 7.7.
If $\dim H_{\mathbb{R}}\geq 3$ and $(U_{t})_{t}$ nontrivial, then $(\Gamma(H_{\mathbb{R}},U_{t}),\phi_{U})$ is selfless, simple, and purely infinite.
Proof.
Since $\phi_{U}$ is faithful and non-tracial, simplicity and pure infiniteness are implied by selflessness [50, Theorem 3.1], so we must show that $\Gamma(H_{\mathbb{R}},U_{t})$ is selfless.
We may split $H_{{\mathbb{R}}}$ as a direct sum of $(U_{t})_{t}$-invariant subspaces: $H_{\mathbb{R}}=H_{c}\oplus H_{wm}$, where $H_{c}$ is a (potentially infinite) direct sum of finite-dimensional subrepresentations, and $H_{wm}$ has no nonzero finite-dimensional subrepresentations. Since $\Gamma(H_{{\mathbb{R}}},U_{t})=\Gamma(H_{c},U_{t})*\Gamma(H_{wm},U_{t})$ (by [55, Theorem 2.11]), and the reduced free product of a selfless $C^{*}$-algebra with any $C^{*}$-algebra is selfless, it suffices to show that one of $\Gamma(H_{c},U_{t})$, $\Gamma(H_{wm},U_{t})$ is selfless. Since $\dim_{{\mathbb{R}}}(H_{{\mathbb{R}}})\geq 3$, we either have that $\dim_{{\mathbb{R}}}(H_{wm})>0$ or $\dim_{{\mathbb{R}}}(H_{c})\geq 3$.
[…]
Since $H_{wm}$ has no finite-dimensional subrepresentations, the above paragraph implies that $E$ has no atoms. Let $p_{t}=E((-\infty,t])$. Since $E$ has no atoms, $t\mapsto p_{t}$ is an SOT-continuous path of projections with $\lim_{t\to-\infty}p_{t}=0$ and $\lim_{t\to\infty}p_{t}=1$. Thus we may find an increasing sequence $0<t_{1}<t_{2}<\cdots$ of real numbers so that $t_{n}\to\infty$ and $p_{t_{i+1}}-p_{t_{i}}\neq 0$. Set $t_{0}=0$, and $K_{i}=(E((-t_{i},-t_{i-1}]\cup[t_{i-1},t_{i})))(K)$ for $i=1,2,\ldots$. By the above paragraph, there are $U_{t}$-invariant subspaces $H_{i}$ of $H_{{\mathbb{R}}}$ with $K_{i}$ the complexification of $H_{i}$ and each $H_{i}\neq 0$. Since $H_{wm}=\bigoplus_{i=1}^{\infty}H_{i}$, we have $\Gamma(H_{wm},U_{t})=*_{i=1}^{\infty}\Gamma(H_{i},U_{t})$. Since $H_{i}\neq 0$, each $\Gamma(H_{i},U_{t})$ has a state-preserving copy of $(C([-2,2]),\mu_{sc})$, where $\mu_{sc}$ is the semicircular measure. By the argument in Example 4 we know that $(C([-2,2]),\mu_{sc})$ is isomorphic to $(C([0,1]),m)$ as $C^{*}$-probability spaces, where $m$ is Lebesgue measure. Thus $(C([-2,2]),\mu_{sc})$ has a state zero unitary. We now apply [50, Theorem 2.8] to see that $\Gamma(H_{wm},U_{t})$ is selfless.
Now suppose $\dim_{{\mathbb{R}}}(H_{c})\geq 3$. If $\dim H_{c}=\infty$, then writing $H_{c}$ as a direct sum of finite-dimensional subrepresentations, we can write $\Gamma(H_{c},U_{t})$ as a direct limit of C*-subalgebras of the form $\Gamma(H_{\mathbb{R}}^{\prime},U^{\prime}_{t})$, with $3\leq\dim H_{\mathbb{R}}^{\prime}<\infty$. By the preservation of selflessness under direct limits [50, Theorem 4.1], it suffices to show that each $\Gamma(H_{{\mathbb{R}}}^{\prime},U_{t}^{\prime})$ is selfless. We may thus assume that $H_{{\mathbb{R}}}$ is finite dimensional.
Suppose that $3\leq\dim H_{\mathbb{R}}<\infty$. In this case $H_{\mathbb{R}}$ decomposes into a direct sum of 1-dimensional or 2-dimensional $U_{t}$-invariant subspaces. By [55, Theorem 2.11], $\Gamma(H_{\mathbb{R}},U_{t})$ can thus be expressed as a reduced free product where each factor is either $\Gamma({\mathbb{R}},\mathrm{id}_{t})$ or $\Gamma({\mathbb{R}}^{2},U_{t})$, with $(U_{t})_{t}$ non-trivial (see [55, Remark 2.12]). We note that $(\Gamma({\mathbb{R}},\mathrm{id}_{t}),\tau_{U})\cong(C[-2,2],\tau)$, with $\tau$ coming from the semicircular distribution. On the other hand, if $(U_{t})_{t}$ acts nontrivially on ${\mathbb{R}}^{2}$, then, by [55, Corollary 4.9],
$(\Gamma({\mathbb{R}}^{2},U_{t}),\phi_{U})\cong(\mathcal{T},\phi_{\lambda})*(C[-2,2],\tau).$
(Here $\mathcal{T}$ denotes the Toeplitz $C^{*}$-algebra and $\phi_{\lambda}$ a certain faithful state on $\mathcal{T}$.) We deduce from these calculations that if $\dim H_{\mathbb{R}}\geq 3$, then
$(\Gamma(H_{\mathbb{R}},U_{t}),\phi_{U})\cong(A,\rho)*(C[-2,2],\tau)*(C[-2,2],\tau),$
for some $(A,\rho)$. It follows that $(\Gamma({\mathbb{R}}^{2},U_{t}),\phi_{U})$ is selfless by Theorem 7.6 and the preservation of selflessness under reduced free products [50, Theorem 4.2]. ∎
@@ 8. Selflessness for free products of finite dimensional abelian $C^{*}$-algebras
We derive this theorem from Theorem 7.1 combined with Dykema’s results on the structure of reduced free products of finite dimensional abelian $C^{*}$-algebras, as well as on their von Neumann algebra counterpart.
@@
Let $p_{2}$ be the unit of $A_{2}$. Since $A$ is simple, $p_{2}$ is a full projection, and so $A$ is stably isomorphic to $p_{2}Ap_{2}$. In the tracial case, selflessness is preserved under stable isomorphism, by [50, Theorems 4.3, 4.4]. It thus suffices to show that $(p_{2}Ap_{2},\frac{1}{\tau(p_{2})}\tau)$ is selfless. By [18, Proposition 2.8],
@@
The conditions (8.1) then read as $\alpha_{i}+\beta_{j}<1$ for all $i,j$. By [21, Theorem 1], these conditions, together with $m+n\geq 5$, are necessary and sufficient for the reduced free product $C^{*}$-algebra $C$ to be a simple $C^{*}$-algebra. They are thus also necessary for $(C,\tau)$ to be selfless.
Let us prove sufficiency. Suppose that $m+n\geq 5$ and that $\alpha_{i}+\beta_{j}<1$ for all $i,j$. Assume without loss of generality that $m\geq 3$ and $n\geq 2$. Let $p=p_{1}+p_{2}$. Since $C$ is a simple $C^{*}$-algebra (by [21, Theorem 1]), $pCp$ is stably isomorphic to $A$. By the permanence of the selfless property under stable isomorphism in the tracial case ([50, Theorems 4.3, 4.4]), it suffices to show that $(pCp,\frac{1}{\tau(p)}\tau)$ is selfless. By [18, Proposition 2.8], we have that
@@
Let us consider first the case when $n+m>5$. In this case, by Dykema’s structure theorem on reduced free products of finite abelian $C^{*}$-algebras [18, Theorem 1],
@@
If $\alpha_{1}+\alpha_{2}+\beta_{j}=1$ for some $j$, then $\alpha_{3}=\beta_{j}$. But this implies that $\alpha_{3}+(1-\beta_{j})=1$, which violates (8.1). Thus, $\alpha_{1}+\alpha_{2}+\beta_{j}\neq 1$ for $j=0,1$. It follows by the structure theorem for reduced free products of the form ${\mathbb{C}}^{2}*{\mathbb{C}}^{2}$ [18, Proposition 2.7], that